OpenStudy (anonymous):

What is the solution to the equation? –4q + 8 = –2q – 16 A. q = –12 B. q = –4 C. q = 4 D. q = 12

OpenStudy (anonymous):

@mathivh

OpenStudy (anonymous):

@alesia/bailey

OpenStudy (anonymous):

@Compassionate @confluxepic

OpenStudy (anonymous):

@hartnn

HanAkoSolo (jamierox4ev3r):

just subtract the like terms for each other? -4q+8=-2q+6 +4q +4q ----------- 8=2q+6 like so. and do the same thing with the 6 and 8

OpenStudy (bringmethehorizon_):

It will be negative i think.

OpenStudy (anonymous):

what 8 and 6?

HanAkoSolo (jamierox4ev3r):

you would subtract 8=2q+6 -6 -6 ------- 2=2q the answer will not be negative @TheAsker2002 does this make sense so far? you just subtract/add the like terms from both sides of the equation

OpenStudy (anonymous):

No :(

OpenStudy (anonymous):

8-6 is 2

HanAkoSolo (jamierox4ev3r):

oh okay. So you see -4q and -2q? these are like terms. Therefore, in an equation, you can treat them as like numbers, just don't forget to include the variable

OpenStudy (bringmethehorizon_):

Can you help me with one too? @Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

and yes, 8-6 is 2 but to get there, you do this first: -4q+8=-2q+6 +4q +4q ----------- 8=2q+6 ^^^ The reason why i was able to add 4q to both sides is because the terms are similar

OpenStudy (alesia/bailey):

–4q + 8 = –2q – 16 okay, so you are trying to isolate q -4q + 8 = -2q - 16 +2q +2q ------> -4q+2q= -2q ------------------- -2q + 8 = - 16 -8 - 8 ------> -16-8= -24 ----------------- -2q = - 24 now you have to get q by itself so you do the opposite and divide -2 on both sides -2q = - 24 ----> -2/-2= 1 and -24/ -2 = 12 1q (or just q) = 12 q=12